OpenStudy (anonymous):

a person expended 500 newtons to move a full wheelbarrow 30 meters how much work was done? a person expended 500 newtons to move a full wheelbarrow 30 meters how much work was done? @Physics

OpenStudy (anonymous):

W = F * D = 500N * 30m = 15000J

OpenStudy (anonymous):

k